OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 3x + 7y = 11? (1, –2) (0,–4) (6, –1) (0, 4)

OpenStudy (mathstudent55):

Try each ordered pair in the equation. If it makes the equation true, that is the one.

OpenStudy (mathstudent55):

To try an ordered pair, take the equation and replace x with the first number and y with the second number. Then do the math and see if that ordered pair makes the equation true.

OpenStudy (mathstudent55):

If you were testing the ordered pair (2, 4), you'd do this: 3x + 7y = 11 3(2) + 7(4) = 11 6 + 28 = 11 34 = 11 Since 34 = 11 is a false statement, (2, 4) is not a point on the graph of the given equation. Now do the same for each choice you have.